# Dynamical Phase Transitions, Caustics, and Quantum Dark Bands

### v1

@@ II Main results: Dynamical Phase Transitions and Caustics
The dynamics in Fock space are determined by the set of quantum amplitudes $\{\langle j,-j|\psi(t)\rangle,\langle j,-j+1|\psi(t)\rangle,\ldots,$ $\langle j,j|\psi(t)\rangle\}$ given by the overlaps between the Fock states and the time-dependent wave function. The results for our example are displayed in Fig. 2. The most striking feature is that there are quite sharply defined boundaries between regions where the amplitudes are finite and where they appear to be zero (more precisely, they are exponentially suppressed as these are the classically forbidden regions, see Sec. IV). These boundaries are fold caustics and are particularly ‘bright’, i.e. regions of high probability. The fold curves meet in pairs at cusp caustics which are the brightest regions of all. Caustics are associated with a breakdown of the corresponding classical ray theory, which will be introduced in detail later, but in the quantum theory they are smooth and dressed by characteristic oscillatory interference patterns which are given by diffraction integrals [75]. Caustics can also be seen in the Loschmidt echo. For the given initial condition the Loschmidt echo, displayed in the right panel of Fig. 2, is also a Fock-state overlap, and hence a cut through the more general Fock-space wave function shown in the left panel of Fig. 2. In a finite system the Loschmidt echo is an analytical function of time. We can identify oscillatory and exponentially decaying regions which are separated by caustics. The exponentially supressed regions are of particular interest here because it is precisely where we find a nonanalytic limit as $N\rightarrow\infty$, and hence a dynamical phase transition, as in Fig. 1.
@@
To build up some intuition for the physics of Fig. 2, we highlight an analogy to a very different physical system: Optical rainbows. Rainbows are caustics in the angular distribution of sunlight scattered by raindrops. In Fig. 3 (left) a photograph of a rainbow is displayed and we can identify three different regions in the sky from bottom to top: A bright region below the primary rainbow, a dark region between the primary and the secondary rainbows (the secondary rainbow is only faintly visible here) and finally another slightly brighter region above the secondary bow. Historically, Airy’s 1838 theory of the rainbow was the first treatment of the wave theory of light near a caustic that correctly tamed the geometric ray singularity to obtain a finite result and was the origin of the eponymous Airy function [79, 80]. The intensity of scattered light as a function of the viewing angle according to this theory is displayed schematically in Fig. 3 (right). It features two back-to-back Airy functions, one for each of the two rainbows. Each rainbow features oscillations (supernumary arcs) on the bright side of the main lobe (the caustic) and exponential decay on the other. The dark region between the primary and secondary rainbows corresponds to Alexander’s DB, after Alexander of Aphrodisias who described it around A.D. 200 [77], and it is only populated by the two exponentially small tails.
The field amplitude in the vicinity of the primary rainbow can be written as [80, 76]
@@
where $k$ is the wavenumber of the incident light, $D_{1}$ is a parameter with units of $\mathrm{length}^{2/3}$ determined by the physical properties of the raindrop, Ai is the Airy function and $\theta_{1}$ is the rainbow angle (about $42^{\circ}$). The same approximation can be made for the secondary rainbow
@@
$I(\theta)\approx|\tilde{C}_{1}(\theta)\exp(-2/3D_{1}^{3/2}k(\theta-\theta_{1})^{3/2})+\tilde{C}_{2}(\theta)\exp(-2/3D_{2}^{3/2}k(\theta_{2}-\theta)^{3/2})|^{2}$ (8)
@@
$r(\theta)=-\frac{1}{k}\log I(\theta)\rightarrow\min\{2/3D_{1}^{3/2}k(\theta-\theta_{1})^{3/2},2/3D_{2}^{3/2}k(\theta_{2}-\theta)^{3/2}\}$ (9)
@@
Figure 3: Left: A rainbow in the skies over Dresden showing Alexander’s DB. The secondary rainbow is only faintly visible. Right: Airy theory of a rainbow for monochromatic light (schematic). One can distinguish two classes of light (red and blue curves) that make up the bright regions. The intensity of light from both the primary and secondary rainbows entering the DB is exponentially suppressed and described by the decaying side of an Airy function. Inside the DB, there exists a critical viewing angle where the two tails have equal magnitude which is the direct analogue of the critical time in a DPT.
@@
$L(t)=|l_{1}(t)+l_{2}(t)|^{2}\,\quad,\qquad l_{i}(t)=\mathrm{e}^{-jr_{i}(t)}$ (10)
where the system size $j$ now takes the role of the wavenumber $k$. Using the same argument as above, the asymptotic rate function will behave nonanalytically as $j\rightarrow\infty$
@@
We can confirm this behavior numerically as displayed in Fig. 4. In the left panel we see that the rate function acquires a finite limit in the DB region, signifying exponential suppression of the return probability. We can identify the typical sharkfin-shape that characterizes DPTs. In the bright regions the return probability has a sub-exponential scaling and hence the rate function becomes zero asymptotically.
Qualitatively this discussion already gives a clear intuition for why the nonanalytic behavior occurs at a DPT. For the FCTFIM we can even obtain a quantitative description using semiclassical methods. In Section IV we employ WKB theory to relate the return probability to classical trajectories [81, 82, 83, 67, 84, 85], similar to the ray theory of geometric optics. Within semiclassics we can compute the two contributions to the Loschmidt echo in Eq. (10) separately. In this way we obtain the exact limiting curve (the sharkfin), as well as accurate finite size predictions, as displayed in Fig. 4 (right panel).
As presented in Sec. V, the semiclassical description also gives the crucial insight for a proposed measurement scheme that can be used to signal the nonanalytic behavior for small size systems, where overlaps can actually be measured [30]. Moreover, our analysis is not restricted to the return probability, but can be applied to arbitrary Fock state overlaps. We elaborate further on this point in Sec. VI, where we introduce the more general concept of a switching line for the Fock representation of a quantum state.
@@
III Rainbows: from rays to waves
In order to gain a better appreciation of the close analogy between rainbows and DPTs, we now outline the mechanism of rainbow formation more fully. Fig. 5 depicts the geometric ray results for a spherical raindrop. Assuming that the light source is far away in comparison to the radius $R$ of the droplet, an incoming ray can be taken as horizontal and its subsequent trajectory is entirely determined by its impact parameter $h$. After an initial refraction upon entry, the trajectory of the ray involves multiple internal reflections but at any of these the ray may escape. Rays escaping at the first reflection contribute to the forward scattered light, but rays escaping at the second and third reflections give rise to the primary (blue) and secondary (red) rainbows, respectively, as shown in the left panel of Fig. 5. The remaining light continues circulating inside the droplet and gives rise to higher rainbows that are increasingly faint. Typically only the primary and occasionally the secondary rainbows are visible to the naked eye.
@@ III Rainbows: from rays to waves
The natural mathematical language for analyzing caustics is catastrophe theory [66]. This describes bifurcations of gradient maps which are essentially theories based on a minimum principle such as the principle of least action. At a bifurcation the number of solutions changes, which in physical terms corresponds to the coalescence or birth of rays, i.e. a caustic. Catastrophe theory specifies a hierarchy of bifurcations of increasing dimension and complexity where the higher ones generally contain the lower ones. The simplest is the fold where two solutions coalesce as a function of a single parameter $x$. The second is the cusp where two parameters $x$ and $y$ are required to fully explore the catastrophe. The cusp is made up of two fold lines that meet at a cusp point. Moving across either of the fold lines the number of solutions changes by two but at the cusp point itself three solutions coalesce. Both the fold and the cusp appear in the dynamics of the TFIM we consider in this paper. Each catastrophe is associated with a normal form or generating function, which in physical problems corresponds to the mechanical action close to the singularity. For the fold and cusp these are
@@
IV Exact Semiclassical Description
We shall now compute the Loschmidt echo within the semiclassical WKB approximation which is valid for large $j$ [81, 82, 83, 67, 84, 85]. This will enable us to understand the occurrence of the nonanalyticity at the DPT and compute the exact asymptotic curve shown in Fig. 4. We start by expressing the time dependent state vector in the energy basis $H|E_{n}\rangle=E_{n}|E_{n}\rangle=j\varepsilon_{n}|E_{n}\rangle$ as
@@ IV Exact Semiclassical Description
where $c_{n}=\langle E_{n}|j,m_{0}\rangle$. For the Loschmidt echo we require the overlap with the initial state
@@
In contrast to the energy mode sum in Eq. (16), this alternative sum converges quickly as long as we do not require long times (in comparison to the period of the classical oscillations).
@@
In our model for any time $t$ there exist different regimes where solutions $T(E)=t$ of the stationary phase condition do or do not exist. This can be seen in Fig. 7, where we display the $T(E)$ curves in an energy-time plot. For short times there either exist two solutions or none. The striking resemblance between this plot and the rainbow case shown in the right panel of Fig. 5 is clearly evident. At either boundary between the classically allowed and forbidden regions two stationary phase solutions coalesce signalling a breakdown of the stationary phase approximation. For the FCTFIM they are the points where fold caustics occur in Fock space. According to our earlier discussion for rainbows, the behavior of the Loschmidt echo can then locally be described in terms of Airy functions. In recogniton of the close connection to rainbow optics, we call the region between the two caustics where no stationary phase solution exists a quantum dark band. For times within the quantum DB the Loschmidt echo is exponentially small, as we observed previously (Fig. 4).
@@
$E_{1/2}$ are the energies of the two returning classical trajectories, i.e. the energies for which $\frac{\mathrm{d}}{\mathrm{d}(E/j)}S_{0}=t$, and $c_{1/2}$ are nonexponential prefactors that we do not explicitly compute. Interference of the two contributions leads to the oscillatory behavior observed in the region to the left of the caustic, as seen in Fig. 4.
@@
The behavior of these saddle point solutions is displayed in Fig. 7. Two real saddle points coalesce at the caustic and then turn into a conjugate pair of complex saddle points. Only one of these points corresponds to an exponentially decaying solution and contributes to the Loschmidt echo in the quantum DB region. From carrying out the steepest descent integration one finds that, to exponential accuracy, the Loschmidt echo is determined by $\tilde{S}_{0}$ evaluated at the saddle point
@@
V A measurement scheme
The Loschmidt echo is an exponentially small quantity that is exponentially hard to measure for a large system, see the definition given in Eq. (1). This makes it impossible to directly observe a dynamical phase transition, which strictly speaking only occurs at infinite system size. It is amusing to contrast this with Alexander’s DB where in a single glance we are effectively observing the results of measurements on a huge number of different systems run for different ‘times’ (angles). A light meter with a high angular resolution could perhaps pick out the angle where the switch between the dominance of the two evanescent tails occurs, although this is made difficult by the exponentially brighter regions above and below the DB.
@@ V A measurement scheme
We expect that this system has the same classical limit as the original FCTFIM. Now let us focus on the nonunitary dynamics of one subsystem described by
$\rho(t)=\mathrm{tr}_{S}|\Psi(t)\rangle\!\langle\Psi(t)|,\qquad\rho(0)=|\psi_{0}\rangle\!\langle\psi_{0}|\,.$ (34)
Because the initial state is pure and separable with respect to the subsystems $I$ and $S$, the mixed state Loschmidt echo of the subsystem $I$ is uniquely defined as [57]
$L(t)=\langle\psi_{0}|\rho(t)|\psi_{0}\rangle,\qquad r(t)=-\frac{1}{N}\ln L(t)\,.$ (35)
This function shows a DPT as in the original model, see Fig. 8. Crucially, one can now perform measurements on the second subsystem in order to extract the $k=0$ and $k=1$ contributions to the Loschmidt echo which make up the two sides of the sharkfin curve in Fig. 8. In particular, we can think of measuring in a single experimental run the Loschmidt echo of subsystem 1 and the $S_{y}$-spin orientation on subsystem 2, because these measurements commute, given that they act on different Hilbert spaces. Then we can post-select the Loschmidt echo measurements depending on the outcomes of the $S_{y}$ measurements. In particular, we form two groups, a ’$+$’ and a ’--’ group, depending on whether the $S_{y}$ outcome was $S_{y}\geq 0$ or $S_{y}<0$. In a semiclassical picture, these two groups isolate the parts of the wave function on the corresponding two halves of the phase space. To formalize this, we consider a measurement of the POVM $\{E_{+},E_{-}\}$ where $\mathbb{1}=E_{+}+E_{-}$ and (considering $N$ even for simplicity)
@@
$\rho_{\pm}(t)=\mathrm{tr}_{S}E_{\pm}|\Psi(t)\rangle\!\langle\Psi(t)|.$ (38)
@@
$L(t)=L_{+}(t)+L_{-}(t),\qquad L_{\pm}(t)=\langle\psi_{0}|\rho_{\pm}(t)|\psi_{0}\rangle\,.$ (39)
Fig. 8 shows the numerical results for the two contributions in a system with just $N=20$ atoms. The two curves clearly make up the two sides of the sharkfin, and cross at the DPT. Because both contributions are exponentially small it is obvious that for $N\rightarrow\infty$ the kink becomes sharp and the curve nonanalytic
@@
VI Switching Lines in Fock Space

### v2

@@ III Caustics in Fock space
given by the overlaps between the Fock states and the time-dependent wave function. The results for our example system (see Sec. II) are displayed in Fig. 2. The most striking feature is that there are quite sharply defined boundaries between regions where the amplitudes are finite and where they appear to be zero (more precisely, they are exponentially suppressed as these are the classically forbidden regions, see Sec. VI). The boundaries are the fold caustics and are particularly ‘bright’, i.e. regions of high probability. The fold curves meet in pairs at cusps which are the brightest regions of all. Caustics can also be seen in the Loschmidt echo. For the given initial condition the Loschmidt echo, displayed in the right panel of Fig. 2, is also a Fock-state overlap, and hence a cut through the more general Fock-space wave function shown in the left panel of Fig. 2. In a finite system the Loschmidt echo is an analytical function of time. We can identify oscillatory and exponentially decaying regions which are separated by caustics. The exponentially supressed ‘dark’ regions are of particular interest here because it is precisely where we find a nonanalytic limit as $N\rightarrow\infty$, and hence a dynamical phase transition, similar to Fig. 1.
IV Main results: Rainbows and Dynamical Phase Transitions
@@
To build up some intuition for the physics of Fig. 2, we highlight an analogy to a very different physical system: Optical rainbows. Rainbows are caustics in the angular distribution of sunlight scattered by raindrops. In Fig. 3 (upper panel) photographs of rainbows are displayed and we can identify three different regions in the sky from bottom to top: A bright region below the primary rainbow, a dark region between the primary and the secondary rainbow (the secondary rainbow is only faintly visible here) and finally another slightly brighter region above the secondary bow. Historically, Airy’s 1838 theory of the rainbow was the first treatment of the wave theory of light near a caustic that correctly tamed the geometric ray singularity to obtain a finite result and was the origin of the eponymous Airy function [80, 81]. The intensity of scattered light as a function of the viewing angle according to this theory is displayed schematically in Fig. 3 (lower panel). It features two back-to-back Airy functions, one for each of the two rainbows. Each rainbow features interference fringes (supernumary arcs) on the bright side of the main lobe (the caustic) and exponential decay on the other. The dark region between the primary and secondary rainbows corresponds to Alexander’s DB, and it is only populated by the two exponentially small tails.
Figure 3: Upper: Examples of rainbows in the skies over Hamilton (left) and Dresden (right) showing primary and secondary bows as well as Alexander’s DB. Lower: Airy theory of a rainbow for monochromatic light (schematic). One can distinguish two classes of light (red and blue curves) that make up the bright regions. The intensity of light from both the primary and secondary rainbows entering the DB is exponentially suppressed and described by the decaying side of an Airy function. Inside the DB, there exists a critical viewing angle where the two tails have equal magnitude which is the direct analogue of the critical time in a DPT.
The field amplitude in the vicinity of the primary rainbow can be written as [81, 82, 78]
@@
where $k=2\pi/\lambda$ is the wavenumber of the incident light, $D_{1}$ is a parameter with units of $\mathrm{length}^{2/3}$ determined by the physical properties of the raindrop, Ai is the Airy function and $\theta_{1}$ is the rainbow angle (about $42^{\circ}$). The prefactor $C_{1}(\theta)$ is non-exponential in $k$ and will not be important in the following. The $2/3$ scaling exponent on $k$ is the Berry index for the fold catastrophe and determines how the fringe spacing depends on the wavenumber of the light. The same approximation can be made for the secondary rainbow
@@
$I(\theta)\approx\left|\tilde{C}_{1}(\theta)\exp\left(-\frac{2k}{3}D_{1}^{3/2}\,(\theta-\theta_{1})^{3/2}\right)+\tilde{C}_{2}(\theta)\exp\left(-\frac{2k}{3}D_{2}^{3/2}\,(\theta_{2}-\theta)^{3/2}\right)\right|^{2}$ (9)
@@
$r(\theta)=-\frac{1}{k}\log I(\theta)\rightarrow\min\left\{\frac{2}{3}D_{1}^{3/2}\,(\theta-\theta_{1})^{3/2},\frac{2}{3}D_{2}^{3/2}\,(\theta_{2}-\theta)^{3/2}\right\}$ (10)
@@
$L(t)=|l_{1}(t)+l_{2}(t)|^{2}\,\quad,\qquad l_{i}(t)=\mathrm{e}^{-jr_{i}(t)}.$ (11)
The system size $j$ now takes the role of the wavenumber $k$ in Eq. (9). This is because in the FCTFIM the classical (ray) limit is equivalent to the thermodynamic limit ($j$ takes the role of 1/$1/\hbar$), in the same way that the ray optics limit corresponds to large wavenumbers $k$. Using the same argument as above, the asymptotic rate function will behave nonanalytically as $j\rightarrow\infty$
@@
We can confirm this behaviour numerically as displayed in Fig. 4. In the left panel we see that the rate function acquires a finite limit in the DB region, signifying exponential suppression of the return probability. We can identify the typical sharkfin-shape that characterizes DPTs. In the bright regions the return probability has a sub-exponential scaling and hence the rate function becomes zero asymptotically.
Qualitatively this discussion already gives an intuitive explanation for the nonanalytic behaviour at a DPT in terms of the caustics forming in the Fock space wave function. For the FCTFIM we can even obtain a quantitative description using semiclassical methods. In Section VI we employ WKB theory to relate the return probability to classical trajectories [83, 84, 85, 73, 86, 87], similar to the ray theory of geometric optics. Within semiclassics we can compute the two contributions to the Loschmidt echo in Eq. (11) separately. In this way we obtain the exact limiting curve (the sharkfin), as well as accurate finite size predictions, as displayed in Fig. 4 (right panel).
As presented in Sec. VII, the semiclassical description also gives the crucial insight for a proposed measurement scheme that can be used to signal the nonanalytic behaviour for small size systems, where overlaps can actually be measured [38]. Moreover, our analysis is not restricted to the return probability, but can be applied to arbitrary Fock state overlaps. We elaborate further on this point in Sec. VIII, where we introduce the more general concept of a switching line for the Fock representation of a quantum state.
@@
V Rainbows: from rays to waves
In order to gain a better appreciation of the close analogy between rainbows and the Fock space caustics that lead to the DPTs in the FCTFIM, we now outline the mechanism of rainbow formation more fully. Fig. 5 depicts the geometric ray results for a spherical raindrop. Assuming that the light source is far away in comparison to the radius $R$ of the droplet, an incoming ray can be taken as horizontal and its subsequent trajectory is entirely determined by its impact parameter $h$, which is defined as the distance of the incoming ray to the symmetry axis of the raindrop (see Fig. 5 left panel). After an initial refraction upon entry, the trajectory of the ray involves multiple internal reflections but at any of these the ray may escape. Rays escaping at the first reflection contribute to the forward scattered light, but rays escaping at the second and third reflections give rise to the primary (blue) and secondary (red) rainbows, respectively, as shown in the left panel of Fig. 5. The remaining light continues circulating inside the droplet and gives rise to higher rainbows that are increasingly faint. Typically only the primary and occasionally the secondary rainbows are visible to the naked eye.
@@ V Rainbows: from rays to waves
Let us now consider caustics from the point of view of catastrophe theory [47]. This describes bifurcations of gradient maps which are essentially theories based on a minimum principle such as the principle of least action. At a bifurcation the number of solutions changes, which in physical terms corresponds to the coalescence or birth of rays, i.e. a caustic. Catastrophe theory specifies a hierarchy of bifurcations of increasing dimension and complexity where the higher ones generally contain the lower ones. The simplest is the fold where two solutions coalesce as a function of a single parameter $x$. The second is the cusp where two parameters $x$ and $y$ are required to fully explore the catastrophe. Moving across either of the two fold lines that make up the cusp the number of solutions changes by two but at the cusp point itself three solutions coalesce. Both the fold and the cusp appear in the dynamics of the TFIM we consider in this paper. Each catastrophe is associated with a normal form or generating function, which in physical problems corresponds to the mechanical action close to the singularity. For the fold and cusp these are
@@
VI Exact Semiclassical Description
We shall now compute the Loschmidt echo for the FCTFIM within the semiclassical WKB approximation which is valid for large $j$ [83, 84, 85, 73, 86, 87]. This will enable us to understand the occurrence of the nonanalyticity at the DPT and compute the exact asymptotic curve shown in Fig. 4. Moreover, the theory establishes a rigorous connection to classical mechanics, similar to the ray optics framework that we discussed for the optical rainbow in the previous section.
We start by expressing the time dependent state vector in the energy basis $H|E_{n}\rangle=E_{n}|E_{n}\rangle=j\varepsilon_{n}|E_{n}\rangle$ as
@@ VI Exact Semiclassical Description
where $c_{n}=\langle E_{n}|j,m_{0}\rangle$. For the Loschmidt echo we need to compute the overlap with the initial state
@@
In contrast to the energy mode sum in Eq. (17), this alternative sum converges quickly as long as we do not require long times (in comparison to the period of the classical oscillations). If the limit of large system size is taken as below, the spreading of the wave function in Fock space will become slow and a semiclassical evaluation of the sum can also be performed for long times, with only a few relevant terms that contribute significantly at each time.
@@
In our model for any time $t$ there exist different regimes where solutions $T(E)=t$ of the stationary phase condition do or do not exist. This can be seen in Fig. 7, where we display the $T(E)$ curves in an energy-time plot. For short times there either exist two solutions or none. The striking resemblance between this plot and the rainbow case shown in the right panel of Fig. 5 is clearly evident. At either boundary between the classically allowed and forbidden regions two stationary phase solutions coalesce signalling a breakdown of the stationary phase approximation. For the FCTFIM they are the points where fold caustics occur in Fock space. According to our earlier discussion for rainbows, the behaviour of the Loschmidt echo can then locally be described in terms of Airy functions. In recogniton of the close connection to rainbow optics, we call the region between the two caustics where no stationary phase solution exists a quantum dark band. For times within the quantum DB the Loschmidt echo is exponentially small, as we observed previously (Fig. 4).
@@
$E_{1/2}$ are the energies of the two returning classical trajectories, i.e. the energies for which $\frac{\mathrm{d}}{\mathrm{d}(E/j)}S_{0}=t$, and $c_{1/2}$ are nonexponential prefactors that we do not explicitly compute. Interference of the two contributions leads to the oscillatory behaviour observed in the region to the left of the caustic, as seen in Fig. 4.
@@
The behaviour of these saddle point solutions is displayed in Fig. 7. Two real saddle points coalesce at the caustic and then turn into a conjugate pair of complex saddle points. Only one of these points corresponds to an exponentially decaying solution and contributes to the Loschmidt echo in the quantum DB region. From carrying out the steepest descent integration one finds that, to exponential accuracy, the Loschmidt echo is determined by $\tilde{S}_{0}$ evaluated at the saddle point
@@
VII A measurement scheme
The Loschmidt echo is an exponentially small quantity that is exponentially hard to measure for a large system, see the definition given in Eq. (1). This makes it impossible to directly observe a dynamical phase transition, which strictly speaking only occurs at infinite system size. It is interesting to contrast this with Alexander’s DB where in a single glance we are effectively observing the results of measurements on a huge number of different systems run for different ‘times’ (angles). A light meter with a high angular resolution could perhaps pick out the angle where the switch between the dominance of the two evanescent tails occurs, although this is made difficult by the exponentially brighter regions above and below the DB.
@@ VII A measurement scheme
We expect that this system has the same classical limit as the original FCTFIM. Now let us focus on the nonunitary dynamics of one subsystem, for instance subsystem $I$, described by the reduced state $\rho_{I}(t)$
$\rho_{I}(t)=\mathrm{tr}_{S}|\Psi(t)\rangle\!\langle\Psi(t)|,\qquad\rho_{I}(0)=|\psi_{0}\rangle\!\langle\psi_{0}|\,.$ (35)
Because the initial state $\rho_{I}(0)$ is chosen as a pure state, the mixed state Loschmidt echo of the subsystem $I$ is uniquely defined as [62]
$L(t)=\langle\psi_{0}|\rho_{I}(t)|\psi_{0}\rangle,\qquad r(t)=-\frac{1}{N}\ln L(t)\,.$ (36)
This function shows a DPT as in the original model, see Fig. 8. Crucially, one can now perform measurements on the second subsystem (subsystem $S$) in order to extract the $k=0$ and $k=1$ contributions to the Loschmidt echo which make up the two sides of the sharkfin curve in Fig. 8. In particular, we can think of measuring in a single experimental run the Loschmidt echo of subsystem $I$ and the $S_{y}$-spin orientation on subsystem $S$, because these measurements commute, given that they act on different Hilbert spaces. Then we can post-select the Loschmidt echo measurements depending on the outcomes of the $S_{y}$ measurements. In particular, we form two groups, a ’$+$’ and a ’--’ group, depending on whether the $S_{y}$ outcome was $S_{y}\geq 0$ or $S_{y}<0$. In a semiclassical picture, these two groups isolate the parts of the wave function on the corresponding two halves of the phase space. To formalize this, we consider a measurement of the POVM $\{E_{+},E_{-}\}$ where $\mathbb{1}=E_{+}+E_{-}$ and (considering $N$ even for simplicity)
@@
$\rho_{I}^{\pm}(t)=\frac{\mathrm{tr}_{S}E_{\pm}|\Psi(t)\rangle\!\langle\Psi(t)|}{\langle\Psi(t)|E_{\pm}|\Psi(t)\rangle}.$ (39)
@@
$L(t)=L_{+}(t)+L_{-}(t),\qquad L_{\pm}(t)=p_{\pm}(t)\langle\psi_{0}|\rho_{I}^{\pm}(t)|\psi_{0}\rangle\,,\qquad p_{\pm}(t)=\langle\Psi(t)|E_{\pm}|\Psi(t)\rangle,$ (40)
where $p_{\pm}(t)$ is the probability of a respective measurement outcome of the $E_{\pm}$ measurement at time $t$. Fig. 8 shows the numerical results for the two contributions in a system with just $N=20$ atoms. The two curves clearly make up the two sides of the sharkfin, and cross at the DPT. Because both contributions are exponentially small it is obvious that for $N\rightarrow\infty$ the kink becomes sharp and the curve nonanalytic
@@
VIII Switching Lines in Fock Space